Can someone please help me with this equation?
This equation, 3x - y = 18, crosses the y axis at ??

Found 2 solutions by Alan3354, drk:
Answer by Alan3354(69443)

This equation, 3x - y = 18, crosses the y axis at
--------
It crosses at x = 0
3*0 - y = 18
y = -18
--> (0,-18)

Answer by drk(1908)

To find out where it crosses the y-axis, cover the x information and solve for y. We get
3*0 - y = 18
0 - y = 18
-y =18
y = -18